OpenStudy (idealist10):

Solve y'=(2y^2+(x^2)(e^(-(y/x)^2)))/(2xy) explicitly.

OpenStudy (idealist10):

y=ux u=y/x y'=u+u'x y'=y/x+(xe^(-(y/x)^2))/2y u'x+u=u+x/(2ye^(u^2)) u'x=(1/2)(1/u)(1/e^(u^2)) u'x=1/(2ue^(u^2)) 2ue^(u^2) du=dx/x t=u^2 dt=2u du e^t dt=dx/x e^t=ln abs(x)+C e^(u^2)=ln abs(x)+C u^2=ln(ln abs(x)+C) u=+/-sqrt(ln(ln abs(x)+C)) y=+/-x*sqrt(ln(ln abs(x)+C)) But the answer in the book says y=+/-x*ln(ln abs(x)+C). So which answer is right? My answer or the answer in the book?

OpenStudy (idealist10):

@hartnn

OpenStudy (anonymous):

is this even possible to answer?

OpenStudy (idealist10):

What do you mean?

OpenStudy (cwrw238):

its certainly complex!!

OpenStudy (idealist10):

You know what, it'd be good if the user sithsandgiggles is here.

hartnn (hartnn):

I don't see any error in your work.

OpenStudy (idealist10):

So the answer in the book must be wrong, then. Right?

hartnn (hartnn):

There could be multiple *equivalent* answers for the same question, but i don't see any way for those 2 answers to be equivalent. So yes, we can safely conclude that the answer in the book is incorrect.

OpenStudy (idealist10):

Thank you!

OpenStudy (cwrw238):

theres certainly got to be a sqrt in the answer